Grade 12Wave Motion

What is meant by angle between vectors? What is resultant of vectors?

Kishore Kumar
9 Years agoGrade 12

3 Answers

REENU SIVADARSHINI.M
7 Years ago
ANGLE BETWEEN VECTOR
IF \Theta IS THE ANGLE BETWEEN THE 2  VECTORS A\rightarrow  B\rightarrow  .     
THE MAGNITUDE OF THE RESULTANT  VECTOR  R\rightarrow CAN BE FOUND USING  TRIANGULAR LAW OF ADDITION.     / R\rightarrow/=  [A2+B2+2AB COS\Theta]1/2
VECTOR R MAKES AN ANGLE \alpha WITH VECTOR A
    tan\alpha=B sin
 
 
REENU SIVADARSHINI.M
7 Years ago
sry i posted the answer without completing it 
THE DIRECTION OF RESULTANT VECTOR R
tan\alpha= B \sin \Theta / A+B\cos \Theta   .
\alpha=  tan-1  B\sin \Theta  / A+B\cos \Theta.
 HERE \Theta IS THE ANGLE BETWEEN THEVECTOR A AND B  WHEREAS  \alpha  IS THE ANGLE BETWEEN THE RESULTANT VECTOR AND VECTOR A.